(51)Suppose the GDP per capita of Bolivia was equal to US$650 and its GDP was equal to US$30 billion in 2019, which of the following represented Bolivia’s population in 2019:

(a)10 million

(b)8 million

(c)More information needed

(d)None of the above

(52)Which of the following is not a component of aggregate investment expenditures in a country’s GDP:

(a)Business investment expenditures

(b)Changes in inventories

(c)Durable consumption expenditures

(d)None of the above

(53)The market value of all final goods and services in an economy produced by resources owned by the nationals of that country:

(a)National income

(b)GDP

(c)GDP per Capita

(d)None of the above

(54)Which of the following activities is not considered part of US GDP accounts?

(a)Market value of surgeries

(b)Market value of cocaine

(c)Market value of steel

(d)Market value of perfumes

(55)Suppose the GDP is equal to $1580b, G=$470b, Ig=$260b, X=$8b, M=$5b, what is the value of C?

(a)$1265b

(b)$853b

(c)$847b

(d)None of the above
Answer #1

51. GDP = $ 30,000 millions

GDP per capita = $ 650

Population = GDP / GDP per capita = 30,000/ 650 = 46.15 million

The correct opinion is d. None of the above

52. Business investment expenditure, changes in inventories, and durable consumption expenditure is a part of GDP. Therefore the correct option is d. None of the above.

53. The market value of all the final goods and services produced by the resources owned by the country's nationals is the National income. The correct opinion is a.

54. Cocaine is illegally produced and is not included in the GDP. Therefore the correct option is b.

55. GDP = C + I + G + (X - M)

1580 = C + 470 + 260 + (8-5)

C = $ 847

The correct opinion is c.